12.Two masses are arranged as shown on a ramp that makes an angle of 36° with the horizontal. Mass m1 is twice mass m2.

. What is the minimum coefficient of static friction required between mass and the ramp to ensure

equilibrium?




1
Expert's answer
2021-01-10T18:25:44-0500
"m_1=2m_2=2m\\2mg\sin{36}=mg+\mu (2m)g\cos{36}
\\2\sin{36}=1+2\mu\cos{36}\\\mu=0.109"